A Carnot engine extracts 910J from a 550K reservoir during each cycle and rejects 500J to a cooler reservoir. It operates at 22 cycles per second.

Find the work done during each cycle.

Find its efficiency.

Find the temperature of the cool reservoir.

Find its mechanical power output.
Answer #1

a)910 - 500 = 410 J

b) e = 1 - (Qh/Qc) = 1 - (500/910) = 0.4505

c)500/910 = Tc/Th = Tc/550 ==> Tc = 302.2 K

d)P = W/T = 410/(1/22) = 9020 W